# E-Gen: Leveraging E-Graphs to Improve Continuous Representations of Symbolic Expressions

### v1

@@ Abstract
As vector representations have been pivotal in advancing natural language processing (NLP), some prior research has concentrated on creating embedding techniques for mathematical expressions by leveraging mathematically equivalent expressions. While effective, these methods are limited by the training data. In this work, we propose augmenting prior algorithms with larger synthetic dataset, using a novel e-graph-based generation scheme. This new mathematical dataset generation scheme, E-Gen, improves upon prior dataset-generation schemes that are limited in size and operator types. We use this dataset to compare embedding models trained with two methods: (1) training the model to generate mathematically equivalent expressions, and (2) training the model using contrastive learning to group mathematically equivalent expressions explicitly. We evaluate the embeddings generated by these methods against prior work on both in-distribution and out-of-distribution language processing tasks. Finally, we compare the performance of our embedding scheme against state-of-the-art large language models and demonstrate that embedding-based language processing methods perform better than LLMs on several tasks, demonstrating the necessity of optimizing embedding methods for the mathematical data modality.
@@ 1 Introduction
While large language models (LLMs) (Devlin, 2018; Radford, 2018; Radford et al., 2019; Rae et al., 2021; Chowdhery et al., 2023; Achiam et al., 2023) have demonstrated effectiveness in processing natural language, these methods still perform suboptimally with math-based content which plays an important role across numerous domains (Zanibbi et al., 2024; Rohatgi et al., 2019). For instance, even a state-of-the-art LLM like GPT-4V only reaches 49% accuracy in MathVista (Lu et al., 2023), a comprehensive benchmark to evaluate the mathematical reasoning capabilities of LLMs. In another instance, as tested in Frieder et al. (2024), GPT-4’s performance in computing integration is dominated by specialized embedding-based models (Lample and Charton, 2019; Lample et al., 2022). Therefore, effective approaches to process semantically rich mathematical notation are necessary.
@@
    The embedding models are evaluated on two out-of-distribution (OOD) downstream tasks, to demonstrate models’ generalizability and robustness.
@@ 2 Related Work
In mathematical language processing (MLP) (Meadows and Freitas, 2022), semantic representation shows strong potential across various problems. A representative application is mathematical information retrieval (MIR) (Kristianto et al., 2016; Zanibbi et al., 2016a; Mansouri et al., 2022c), where expressions or keywords are ranked based on their relevance to a query. Gao et al. (2017) proposes Symbol2Vec, a mathematical symbol representation generation scheme to group LaTeX symbols having similar contexts together. They further extend this approach to a MIR scheme (Formula2Vec) to prove its effectiveness. Krstovski and Blei (2018) converts symbolic layout trees (Zanibbi et al., 2016b) of mathematical expressions into token sequences and generates representation based on word2vec. Mansouri et al. (2019) further improves this approach by combining the symbol layout tree and operator tree representations of an expression. Peng et al. (2021) proposes MathBERT, a pretrained model based on MIR and formula headline generation task (Yuan et al., 2020).
@@ 3.1 Corpus Generation
To generate a diverse set of mathematical expressions, we manually design templates containing placeholders for arithmetic operators, functional operators, and numerical values. Expressions are systematically instantiated by replacing the placeholders in templates with all possible operators and values, resulting in approximately 5,000 raw expressions in prefix notation. The templates are carefully designed to cover fundamental arithmetic and functional operators while maximizing the potential for mathematical transformations. Each of the resulting raw expressions is processed by E-Gen, which applies around 800 mathematical rules to generate clusters of semantically equivalent expressions.
@@
Figures 1(a) to 1(d) show the e-graph saturation process. In the initial e-graph (Figure 1(a)), the e-node “$+$” is linked to two child e-classes “$x$” and “$8$”, respectively, and the e-node “--” is linked to a subgraph “$(x+8)$” and a child e-class “$8$”, together forming the raw expression “$(x+8)-8$”. In Figure 1(b), a new e-node “$+$” is added to the top e-class after the associative law is introduced. The two e-nodes, “--” and “$+$” in the top e-class represent subgraphs for expressions “$(x+8)-8$” and “$x+(8-8)$” which are equivalent rewrites generated by the associative law. Similarly, transformations such as “$(x-x)\rightarrow 0$” and “$x+0\rightarrow x$” are embedded into the e-graph in the steps shown in Figures 1(c) and 1(d). The e-graph is iteratively expanded by applying each applicable rule, thereby capturing all possible equivalent expressions.
@@
$[ec_{i}]$ $\rightarrow[var/num]$
$[ec_{i}]$ $\rightarrow[op\;ec_{j}\;ec_{k},\ldots]$
where $ec$ denotes eclasses, $op$ and $var/num$ denotes enodes representing operators and variables/numbers.
@@ 3.2 Corpus Analysis
The new corpus encompasses a comprehensive set of arithmetic and functional operators as shown in Table 1. Table 2 presents partial examples from two clusters of equivalent expressions generated by E-Gen. The corpus is structured into clusters, where each cluster contains numerous mathematically equivalent expressions derived through a series of flexible transformations. These transformations span both simple and complex mathematical relationships, significantly enhancing expression diversity and enabling models to learn underlying mathematical rules more efficiently with fewer raw expressions.
Raw expression: $\tanh(3x-(-4))-6$
@@
Table 2: Examples of equivalent expressions generated with E-Gen. Expressions listed below each raw expression are part of the equivalent rewrites. Additional examples are in Appendix A.2.
As detailed in Table 3, the E-Gen corpus not only supports a wider range of operators, but also achieves a significantly higher average cluster size of 102, compared to an existing corpora Equivalent Expressions Dataset (EED) (Gangwar and Kani, 2023) that is generated using SymPy and limited to a cluster size of 2 (pairs). This substantial increase in cluster size enables models to develop a deeper semantic understanding of the diversity of mathematical transformations.
@@ 5.1 Evaluation Tasks
K-Means clustering (Macqueen, 1967) is employed to evaluate the performance of both seq2seq and contrastive learning models. The test set comprises 8,077 expressions in 279 clusters, with cluster sizes ranging from 20 to 40 expressions. Subsequently, K-Means clustering is applied to group the expressions into clusters. A visualization of clusters is shown in Figure 3. To evaluate clustering performance, the predicted cluster labels from the K-Means algorithm are mapped to the corresponding ground truth labels, with accuracy computed as the proportion of expressions correctly assigned to their respective clusters.
@@
As shown in Table 4, both the seq2seq and contrastive learning models, trained on the E-Gen cluster-based dataset, demonstrate strong performance in grouping equivalent expressions while separating non-equivalent ones. Notably, the CL models slightly outperform the seq2seq model in this task because they are trained to explicitly group similar expressions together. In contrast, SemEmb, which is trained on a SymPy-generated corpus, only achieves 37.70% accuracy. The lower performance is attributed to the limited diversity of mathematical transformations in the SymPy-generated corpus which restricts the model’s ability to effectively capture and distinguish small and large syntactic variations among equivalent expressions.
@@
To evaluate whether models trained on the E-Gen corpus capture mathematical semantics rather than relying on syntactic similarity, we design an experiment where the model must identify the single semantically equivalent expression among syntactically similar candidates. Specifically, we sample 1,000 query expressions from the test set and, for each query, we introduce one correct answer that is semantically equivalent but syntactically distinct from the query. Additionally, six distractors are generated by making minor syntactic modifications over the query and the correct answer (e.g., altering a single operator or numerical value). These distractors are divided into two groups: three closely resembling the query and three closely resembling the correct answer. The model is tasked with selecting the correct equivalent expression from seven candidates. In the example shown in Table 5, the candidate highlighted in bold represents the correct answer, which is syntactically distinct from the query involving the derivative operation. Candidate 4, 5, 7 are syntactically similar to the correct answer, while candidate 1, 3, 6 shares a similar structure with the query. As shown in Table 6, models trained on the E-Gen corpus significantly outperform SemEmb, demonstrating the models’ ability to distinguish between semantically equivalent but syntactically different expressions, as well as syntactically similar yet semantically non-equivalent expressions.
@@
$\frac{d}{dx}3x+\sin(x-4)/6$
$\boldsymbol{2+\cos(x-4)/6}$
$\frac{d}{dx}2x+\sin(x+4)/6$
$62+\cos(x/4)/$ 6
$82+\tan(x-4)/$ 8
$\frac{d}{dx}2x+\cos(x-4)/6$
$69+\cos(x-4)/$ 6
Table 5: Examples of semantic understanding beyond syntactic similarity query, where the correct option is written in red. Six distractors are generated by slightly changing the query expression and the correct option to evaluate model’s capacity in distinguishing syntactically similar but semantically different expressions
@@
Mistake detection in mathematical derivations is a downstream OOD task of which test set is generated by Sympy to test the robustness and generalizability of the models.
Derivation seq2seq GPT-4o
$-\frac{1}{8}(\sinh(\cosh^{-1}(7-\frac{x}{-8})))^{-1}$
$-\frac{1}{8}(\sinh(\cosh^{-1}(7+\frac{x}{8})))^{-1}$
$-\frac{1}{8}(\sqrt{6+\frac{x}{8}}\sqrt{8+\frac{x}{8}})^{-1}$bigstar
$-\frac{1}{8}(6+\frac{x}{8})^{-\frac{1}{2}}(8+\frac{x}{8})^{-\frac{1}{2}}$
$-\frac{1}{8}(6+\frac{x}{8})^{-\frac{1}{2}}(8-7x)^{-\frac{1}{2}}$
$5/\csc(\csc^{-1}(1/\ln(\frac{x}{-5})))$
$5/\csc(\csc^{-1}(1/\ln(\frac{1}{-5}x)))$
$5/{\color[rgb]{1,0,0}\sec}(\csc^{-1}((\ln(\frac{1}{-5}x))^{-1}))$bigstar
$5/(1-(\ln(\frac{1}{-5}x))^{2})^{-\frac{1}{2}}$
$5\sqrt{1-(\ln(\frac{1}{-5}x))^{2}}$
In our work, we first generate a step-by-step mathematical derivation based on expressions in the original test set of the E-Gen corpus. Then, mistakes are introduced to randomly selected steps for each derivation. Models are expected to identify steps with mistakes in the derivation steps. Specifically, transformations between consecutive derivation steps are classified as either “mistake” or “no mistake”. 80% derivations are used to calculate a semantic similarity threshold as the threshold for binary classification [^1] Similarity threshold calculation is shown in Appendix C.2, with the remainder being the test set. A step is flagged as a potential mistake when the cosine similarity between prior steps and the current step is below the threshold. Two examples of mistake detection are shown in Table 8, some steps with minor mistakes often share a similar structure with the preceding steps, such as mistakes steps shown in Table 8, making some incorrect transformations between consecutive steps appear deceptively more similar to correct mathematical transformations that significantly change the expressions’ structure. This poses a significant challenge to models’ ability to understand mathematical semantics.
Model Precision Recall F1
no mistake seq2seq 96.40 94.69 95.54
CL Mean 97.59 92.52 94.99
CL Max 97.78 91.93 89.14
SemEmb 92.92 83.49 87.95
mistake seq2seq 74.68 81.61 77.99
CL Mean 69.33 88.10 77.60
CL Max 67.96 89.14 77.12
SemEmb 44.46 67.50 53.61
The test set has 18,462 derivation steps, including 2,974 steps containing mistakes, with the remainder being error-free. As shown in Table 9, precision, recall, and F1-score are used to evaluate the performance. All three models trained on E-Gen corpus demonstrate strong effectiveness in identifying potential mistakes and better generalizability in OOD task, significantly outperforming the SemEmb approach.
Embedding Algebra.
Embedding algebra is a classic task to evaluate if embeddings capture semantic information of an word/token. Techniques like word2vec (Mikolov, 2013) and GloVe (Pennington et al., 2014) exhibit the ability to perform analogy-based reasoning through algebraic operations on their representation vectors, facilitating solutions to analogies such as “Berlin is to Germany as Paris is to France”. By extending this task to mathematical analogies, we test if models can really understand mathematical transformations or just have a shallow understanding of expressions’ structure. For a given triplet of expressions $\boldsymbol{x}_{1}$, $\boldsymbol{y}_{1}$, and $\boldsymbol{x}_{2}$, we compute:
$f(\boldsymbol{\hat{y}}_{2})=-f(\boldsymbol{x}_{1})+f(\boldsymbol{y}_{1})+f(\boldsymbol{x}_{2})$ (2)
where $f$ denotes a function $f:\mathcal{X}\rightarrow\mathcal{Z}$, which maps an expression $\boldsymbol{x}$ to its representation vector $\boldsymbol{z}$. We then identify the expression whose embedding vector is most similar to $f(\boldsymbol{\hat{y}}_{2})$ in terms of cosine similarity, excluding the original expressions $\boldsymbol{x}_{1}$, $\boldsymbol{y}_{1}$, and $\boldsymbol{x}_{2}$ from consideration.
For this experiment, 584 analogy examples are manually constructed, and all expressions from the E-Gen corpus are used as the search pool, removing those equivalent to $\boldsymbol{x}_{2}$ and $\boldsymbol{y}_{2}$ to ensure uniqueness of answer, while adding any necessary expressions to complete the analogy.
Model Accuracy (%)
seq2seq 70.38
CL Mean 64.73
CL Max 50.34
SemEmb 54.85
GPT-4o 39.60
As shown in Table 10, the seq2seq model achieves the highest accuracy, indicating its ability to effectively learn underlying mathematical rules and handle basic substitutions. Conversely, SemEmb exhibit low accuracy on this task and have a tendency to imitate the transformation between $\boldsymbol{x}_{1}$ and $\boldsymbol{y}_{1}$, leading to “look-alike” predictions rather than a true understanding of the mathematical rules, such as Test 1, 4, 5. Examples are shown in Table 7.
5.2 Comparison with GPT-4o
To further assess the quality of vector representations of symbolic expressions, a comparative analysis is conducted between the seq2seq model, trained on the E-Gen corpus, and the state-of-the-art large language model GPT-4o (Achiam et al., 2023) on several tasks.
Mistake Detection.
For mistake detection, a small test set is randomly sampled from the mistake detection test set mentioned above, consists of 322 derivation steps, of which 50 contain mistakes and the remainder are error-free. We first provide GPT-4o a derivation example in the prompt, asking it to identify potential mistakes. If GPT-4o fails to find the potential mistake, we teach it where the mistake appears until it can understand the task and pass the test. Once verified, we evaluate GPT-4o using derivation queries from the test set. As shown in Table 11, our work outperforms GPT-4o across all metrics. GPT-4o demonstrates a recall of 66%, suggesting its capability of detecting a notable portion of errors. However, its low precision indicates a tendency to misclassify correct transformations as mistakes, leading to a high rate of false positives.
Model Precision Recall F1
no mistake seq2seq 96.98 94.49 95.72
CL Mean 97.64 91.18 94.30
CL Max 97.78 91.93 94.77
GPT-4o 93.58 90.84 92.19
mistake seq2seq 73.68 84.00 78.50
CL Mean 64.71 88.00 74.58
CL Max 67.96 89.14 77.12
GPT-4o 56.90 66.00 61.11
Table 8 illustrates examples of GPT-4o’s misjudgments, which commonly occur during complex transformations, particularly those involving function operators or relatively intricate arithmetic calculations. For instance, in Example 1, GPT-4o performs suboptimally with evaluating equivalence when expressions are syntactically different (e.g., $\sinh(\cosh^{-1}(x))=\sqrt{x+1}\sqrt{x-1}$). In Example 2, GPT-4o fails to identify the operator substitution in step 3, judging it as a valid derivation.
Embedding Algebra.
We also conduct the embedding algebra task on GPT-4o, which performs worse than both the seq2seq and CL embedding models, achieving only 39.60% accuracy on the test set as shown in Table 10. Table 7 provides examples comparing the performance of the seq2seq model and GPT-4o on embedding algebra tests. Similar to the mistake detection test, an example query is presented to GPT-4o to verify its understanding and confirm that it produces the correct output before proceeding with the test. While the seq2seq model consistently makes accurate predictions by adhering to mathematical rules, GPT-4o shows a partial understanding of certain mathematical properties, such as correctly predicting the periodicity of “$\csc(x)$”. However, it often imitates the structure of “$\boldsymbol{y}_{1}$” without fully considering the underlying mathematical rules. For instance, in Test 5, GPT-4o incorrectly imitates the structure of “$\boldsymbol{y}_{1}$”, subtracts “$\pi/2$” from “$x$” in “$\csc(x)$”, while the correct answer should be “$\csc(x+\pi/2)$” which is mathematically equivalent to “$\csc(x)$”. A similar error occurs in Test 8, where GPT-4o predicts “$\boldsymbol{y}_{2}$” as “$\operatorname{csch}^{-1}(x)$” by imitating “$\boldsymbol{y}_{1}:\csc^{-1}(x)$”, instead of applying the correct mathematical transformation. This comparison highlights the effectiveness of E-Gen corpus in helping models to build up understanding in mathematical rules and transformations.
6 Conclusion
In this work, we enhance semantic representations of symbolic expressions by developing E-Gen, a novel mathematical corpus generation framework based on the e-graph data structure. E-Gen shows strong capability in generation of a scalable cluster-based corpus with high diversity in mathematical transformation. To evaluate the effectiveness of the new corpus, two training approaches based on seq2seq and contrastive learning respectively are implemented to capture equivalence relation between expressions. We demonstrate the efficacy of these embedding models for a variety of downstream tasks such as: clustering, mathematical information retrieval, mistake detection, and mathematical analogies. Crucially, we show that these semantic representations outperform prior work and GPT-4o in both quantitative and qualitative tests. This work provides an algorithmic foundation for processing symbolic mathematics and the vector-based representations can integrate easily with vector embeddings of other data modalities.
7 Limitations
This work introduces and evaluates the potential of a novel mathematical corpus generation framework to augment mathematical semantic understanding. While the results are promising, there are several areas for improvement. First, extending the dataset to include a larger variety of operators used in published datasets, such as ArXMLiv (Kohlhase et al., 2024) and ARQMath (Mansouri et al., 2022b), would make the vector embeddings more applicable to real world data.
Following that, more efficient grammar enumeration techniques, such as those based on the Earley algorithm (Earley, 1970), will allow us to more easily extend the E-Gen corpus to include expressions with more complex operators and larger arity. In this manuscript, equivalent expressions are generated by enumerating the grammar obtained from the e-graph using a simple recursive function. Additionally, incorporating variable characteristics, such as dimensionality, phase, and/or bounds, into the embeddings would be useful and an avenue yet to be explored in modern literature.
Another limitation is the scarcity of real-world datasets that focus on symbolic mathematics. While tasks like mathematical information retrieval have been explored in competitions such as NTCIR (Zanibbi et al., 2016a) and ARQMath (Mansouri et al., 2022b), these typically involve a mix of symbolic math and natural language in queries and results. Integration of the our embedding methods with natural language vectors is a research area that has yet to be explored.
Appendix A Cluster-Based Corpus
A.1 Generation Efficiency
Efficiency is a key property of any corpus generation scheme. To evaluate the E-Gen approach, the average time for e-graph saturation and expression rewrite extraction is measured. On average, e-graph saturation process takes 36.83ms, while extraction takes 1.46s per expression. The process is run on an Intel Xeon E5-2666 CPU with a 25-token length limit and a 600s time limit for each raw expression.
A.2 Cluster Example
As discussed in Section 3, the E-Gen framework enables wider coverage of operators, such as derivatives, and the application of a diverse range of mathematical rules. Table 12 shows additional examples of equivalent expression clusters. The key distinction between the E-Gen corpus and prior datasets is that E-Gen consists of clusters of equivalent expressions, rather than simple expression pairs. Each cluster contains a raw expression along with numerous equivalent rewrites. This diverse set of transformations significantly improves the performance of pretrained models in understanding the semantics of mathematical transformations.
Raw expression: $(\sin(2x+5))^{-8}$
$(\sin(5-(-2)\sin(\sin^{-1}(x))))^{-8}$
$(\csc(\pi+(-5-2x)))^{8}$
$(\cos(\pi/2-2x-5)))^{-8}$
$\left(\frac{\cot(2x+5)}{\cos(2x+5)}\right)^{8}$
$(\sin(2(x+\pi)+5))^{-8}$
Raw expression: $\sec^{-1}(7x+6)/(-7)$
$\cos^{-1}(\frac{1}{6-(-7x)})/(-7)$
$(\sec^{-1}(-7x-6)-\pi))\times\frac{1}{7}$
$(\frac{\pi}{2}-\csc^{-1}(7x+6))/(-7)$
$(-1)(\sec^{-1}(7/\sec(\cos^{-1}(x))+6))/7$
$(\cos^{-1}(\cot(\tan^{-1}(7x+6))))/(-7)$
Raw expression: $(\csc(x/7))^{-3}-(-3)$
$31/\left(\csc\left(\frac{\tan(\tan^{-1}x)}{7}\right)\right)^{3}+$ 3
$(\tan(\cot^{-1}(\sin(x/7))))^{-3}+3$
$3+\left(1/\cos(\sec^{-1}(\sin(x/7)))\right)^{3}$
$-1(-3-(\csc(x/7))^{-3})$
$(\sin(x/7))^{3}+3$
$(\sin(\csc^{-1}(\sin(7^{-1}x)))^{-3}-(-3)$
Raw expression: $\frac{d}{dx}(-4x-8)^{-4}+7x$
$(-1)\times(8-(-4x))^{-5}\times 16+7$
$-16\times(4x+8)^{-5}+7$
$7+(-16)\times\left(-1/(-4x-8)^{5}\right)$
$7-(-16)\times(-1)\times(8-(-4x))^{-5}$
$0-(-16(-4x-8)^{-5}-7)$
Raw expression: $\frac{d}{dx}(-9\sinh(-7x+2))$
$63/\operatorname{sech}(-7x+2)$
$-9/(-1/7\cosh(7x-2))$
$63\cosh(2-7x)$
$\frac{63}{2}(\exp(2-7x)+\exp(7x-2))$
$7\times 9/\operatorname{sech}(7x-2)$
Table 12: Additional examples of equivalent expressions generated with E-Gen. Expressions listed below each raw expression are all equivalent to it.
Appendix B Training Details
Both the seq2seq and contrastive learning models are trained using a transformer architecture with 8 attention heads, 6 layers, a model dimensionality of 512, a feedforward dimension of 2048, and a batch size of 256. All models are optimized using the AdamW optimizer (Loshchilov, 2017) with a learning rate of $10^{-4}$ and the CosineAnnealingWarmRestarts scheduler (Loshchilov and Hutter, 2016). Training is conducted on a cluster equipped with an AMD EPYC 7763 64-Core Processor and a 46GB L40S GPU. The training time of each model is shown in Table 13.
Model Training Time (h:min)
seq2seq 38:04
CL Mean 31:56
CL Max 29:59
Table 13: Training time of seq2seq, CL Mean, and CL Max models.
Appendix C Experiments
C.1 K-Means Clustering
Accuracy Calculation.
$\textit{acc}=\frac{1}{K}\sum_{i=1}^{K}\textit{acc}_{i}$ (3)
C.2 Mistake Detection

### v2

@@ Abstract
Vector representations have been pivotal in advancing natural language processing (NLP), with prior research focusing on embedding techniques for mathematical expressions using mathematically equivalent formulations. While effective, these approaches are constrained by the size and diversity of training data. In this work, we address these limitations by introducing E-Gen, a novel e-graph-based dataset generation scheme that synthesizes large and diverse mathematical expression datasets, surpassing prior methods in size and operator variety. Leveraging this dataset, we train embedding models using two strategies: (1) generating mathematically equivalent expressions, and (2) contrastive learning to explicitly group equivalent expressions. We evaluate these embeddings on both in-distribution and out-of-distribution mathematical language processing tasks, comparing them against prior methods. Finally, we demonstrate that our embedding-based approach outperforms state-of-the-art large language models (LLMs) on several tasks, underscoring the necessity of optimizing embedding methods for the mathematical data modality. The source code and datasets are available at https://github.com/MLPgroup/E-Gen.
@@ 1 Introduction
While large language models (LLMs) (Jiang et al., 2023; Dubey et al., 2024; Hurst et al., 2024; Jaech et al., 2024) have demonstrated effectiveness in processing natural language, these methods still perform suboptimally with math-based content which plays an important role across numerous domains (Zanibbi et al., 2024; Rohatgi et al., 2019). For instance, even a state-of-the-art LLM like GPT-4V only reaches 49% accuracy in MathVista (Lu et al., 2023), a comprehensive benchmark to evaluate the mathematical reasoning capabilities of LLMs. In another instance, as tested in Frieder et al. (2024), GPT-4’s performance in computing integration is dominated by specialized embedding-based models (Lample and Charton, 2019; Lample et al., 2022). Therefore, effective approaches to process semantically rich mathematical notation are necessary.
@@
    The embedding models are evaluated on two out-of-distribution downstream tasks, to demonstrate models’ generalizability and robustness.
@@ 2 Related Work
In MLP (Meadows and Freitas, 2022), semantic representation shows strong potential across various problems. A representative application is mathematical information retrieval (MIR) (Kristianto et al., 2016; Zanibbi et al., 2016a; Mansouri et al., 2022c), where expressions or keywords are ranked based on their relevance to a query. Gao et al. (2017) proposes Symbol2Vec, a mathematical symbol representation generation scheme to group LaTeX symbols having similar contexts together. They further extend this approach to a MIR scheme (Formula2Vec) to prove its effectiveness. Krstovski and Blei (2018) converts symbolic layout trees (Zanibbi et al., 2016b) of mathematical expressions into token sequences and generates representation based on word2vec. Mansouri et al. (2019) further improves this approach by combining the symbol layout tree and operator tree representations of an expression. Peng et al. (2021) proposes MathBERT, a pretrained model based on MIR and formula headline generation task (Yuan et al., 2020).
@@ 3.1 Corpus Generation
To generate a diverse set of mathematical expressions, we manually design templates containing placeholders for arithmetic operators, functional operators, and numerical values. Expressions are systematically instantiated by replacing the placeholders in templates with all possible operators and values, resulting in approximately 5,000 initial expressions in prefix notation. The templates are carefully designed to cover fundamental arithmetic and functional operators in various formats while maximizing the potential for mathematical transformations. Each of the resulting initial expressions is processed by E-Gen, which applies around 800 mathematical rules to generate clusters of semantically equivalent expressions.
@@
Figures 1(a) to 1(d) show the e-graph saturation process. In the initial e-graph (Figure 1(a)), the e-node “$+$” is linked to two child e-classes “$x$” and “$8$”, respectively, and the e-node “--” is linked to a subgraph “$(x+8)$” and a child e-class “$8$”, together forming the initial expression “$(x+8)-8$”. In Figure 1(b), a new e-node “$+$” is added to the top e-class after the associative law is introduced. The two e-nodes, “--” and “$+$” in the top e-class represent subgraphs for expressions “$(x+8)-8$” and “$x+(8-8)$” which are equivalent rewrites generated by the associative law. Similarly, transformations such as “$(x-x)\rightarrow 0$” and “$x+0\rightarrow x$” are embedded into the e-graph in the steps shown in Figures 1(c) and 1(d). The e-graph is iteratively expanded by applying each applicable rule, thereby capturing all possible equivalent expressions.
@@
$e\langle\text{index}\rangle$ $\rightarrow[var/num]$
$e\langle\text{index}\rangle$ $\rightarrow[op\;e\langle\text{index}\rangle\;e\langle\text{index}\rangle,\;\ldots]$
where $e\langle\text{index}\rangle$ denotes eclass with corresponding index, $op$ and $var/num$ denote enodes representing operators and variables/numbers.
@@ 3.2 Corpus Analysis
The new corpus encompasses a comprehensive set of arithmetic and functional operators as detailed in Table 1. Table 2 presents examples from three clusters of equivalent expressions generated by E-Gen. The corpus is structured into clusters, each containing numerous mathematically equivalent expressions derived through a series of flexible transformations. These transformations span both simple and complex mathematical relationships, significantly enhancing expression diversity and enabling models to efficiently learn underlying mathematical rules with fewer initial expressions.
Initial expression: $(\frac{x}{6}+2)^{9}+3x$
$-(-2-\frac{x}{6})^{9}+(-3(-x))$
$3x-(-2-\frac{1}{6}x)^{9}$
$(-1)(-3x-(2+\frac{x}{6})^{9})$
$1/(\frac{x}{6}+2)^{-9}+3x$
$(-2+(-1)x/6)^{9}+3x$
Initial expression: $\tanh(3x-(-4))-6$
@@
Table 2: Examples of equivalent expressions generated with E-Gen. Expressions listed below each initial expression are part of the equivalent rewrites. Additional examples are in Appendix A.1.
As detailed in Table 3, the E-Gen corpus supports a wider range of operators and achieves a significantly higher average cluster size of 102, compared to the Equivalent Expressions Dataset (EED) (Gangwar and Kani, 2023) that is generated using SymPy and limited to a cluster size of 2 (pairs). This substantial increase in cluster size enables models to develop a deeper semantic understanding of the diversity of mathematical transformations.
@@ 5.1 Evaluation Tasks
K-Means clustering (Macqueen, 1967) is employed to evaluate the performance of both seq2seq and CL models. The test set comprises 8,077 expressions in 279 clusters, with cluster sizes ranging from 20 to 40 expressions. The K-Means algorithm is applied to group the expressions into clusters. Clustering performance is evaluated by mapping the predicted cluster labels to the corresponding ground truth labels. Accuracy is computed as the proportion of expressions correctly assigned to their respective clusters. A visualization of clusters is shown in Figure 3.
@@
As shown in Table 4, both the seq2seq and CL models trained on the E-Gen cluster-based dataset exhibit strong performance in clustering semantically equivalent expressions while effectively separating non-equivalent ones. The CL models slightly outperform the seq2seq model due to their explicit training objective of grouping semantically equivalent expressions in the latent space. In contrast, SemEmb, trained on a SymPy-generated corpus, achieves only 37.70% accuracy. This significant performance gap underscores the limitation of the SymPy-generated dataset, which lacks the diversity of mathematical transformations present in the E-Gen corpus. Consequently, SemEmb struggles to accurately capture and differentiate both subtle and substantial syntactic variations among equivalent expressions.
@@
To assess whether models trained on the E-Gen corpus capture mathematical semantics rather than relying on syntactic similarity, we design an experiment where the model must identify the single semantically equivalent expression among syntactically similar candidates. Specifically, we sample 1,000 query expressions from the test set. For each query, one correct answer is introduced, which is semantically equivalent but structurally distinct from the query. Additionally, six distractors are generated by making minor syntactic modifications to both the query and the correct answer (e.g., altering a single operator or numerical value). These distractors are divided into two groups: three closely resembling the query and three closely resembling the correct answer. The model is tasked with selecting the correct equivalent expression from seven candidates.
As illustrated in Table 5, the correct answer (bolded) is syntactically distinct from the query. Candidate 2, 4, 7 exhibit structural similarity to the correct answer, while candidate 1, 3, 6 closely resemble the query. As shown in Table 6, models trained on the E-Gen corpus significantly outperform SemEmb, demonstrating their ability to differentiate semantically equivalent expressions despite structural variations, as well as to distinguish between syntactically similar yet semantically non-equivalent expressions.
@@
$1$ $\frac{d}{dx}3x+\sin(x-4)/6$
$2$ $82+\tan(x-4)/$ 8
$3$ $\frac{d}{dx}2x+\sin(x+4)/6$
$4$ $62+\cos(x/4)/$ 6
$5$ $\boldsymbol{2+\cos(x-4)/6}$
$6$ $\frac{d}{dx}2x+\cos(x-4)/6$
$7$ $69+\cos(x-4)/$ 6
Table 5: Examples of semantic understanding beyond syntactic similarity. The correct answer is highlighted in bold. Six distractors are generated by introducing minor modifications to the query and the correct answer to assess the model’s ability to distinguish between syntactically similar but semantically different expressions.
@@
Mistake detection in mathematical derivations is an out-of-distribution downstream task, with the test set generated using SymPy to assess the robustness and generalizability of the models.